OpenStudy (anonymous):

WILL GIVE MEDAL AND WILL FAN solve for x: -2|x - 3| = -12 A. x = -3, x= 4.5 B. x = 3, x = 9 C. x = -3, x = 9 D. no solutions

OpenStudy (anonymous):

C -2(-x+3)=-12 >> -x+3=6 >> x=-3 -2(x-3)=12 >> x-3=-6 >> x=9